OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the given points. (3,5) and (0,4) I don't want the answer i just want someone to tell me how to do this

OpenStudy (anonymous):

Well, slope intercept form is of the notion: y = mx+b where m is the slope and b is the y - intercept. So to figure out the slope you can use this following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] once you have the slop you can start putting it into point slope form which is the following: \[y-y _{1}=m(x-x _{1})\] where (x1,y1) are your first given coordinates, once you plug everything in you can set it into slope - intercept form.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Np :)

OpenStudy (anonymous):

@iambatman would the answer be y= -1/4 x + -2 1/4 ??

OpenStudy (anonymous):

\[m = \frac{ 4-5 }{ 0-3 } = \frac{ -1 }{ -3 } = \frac{ 1 }{ 3 }\] (our slope) \[y-y_1=m(x-x_1) \implies y - 5=\frac{ 1 }{ 3 }(x-3) \implies y -5 = \frac{ 1 }{ 3 }x-0 \implies y = \frac{ 1 }{ 3 }x+5\]

OpenStudy (anonymous):

\[y = \frac{ 1 }{ 3 }x+5\]

OpenStudy (anonymous):

I think I understand what I did wrong but not 100% sure

OpenStudy (anonymous):

That's alright, long as you try is all that matters :) you will understand it!